OpenStudy (anonymous):

MEDAL AND FAN If f(x) is the height, in cm, of a sunflower plant which is x days old, which of the following statements best describes the meaning of f(60) = 210? The height of the sunflower plant is 60 cm when it is 210 days old. The height of the sunflower plant is 210 cm when it is 60 days old. The height of the sunflower plant is 210 cm when it is 3.5 days old. The height of the sunflower plant is 60 cm when it is 3.5 days old.

OpenStudy (math&ing001):

It becomes simple when you identify each variable: x= nbr of days old f(x)= height in cm Could you guess now ?

OpenStudy (anonymous):

I'm just having a hard time because of the wording of the answers I have dyslexia. But my guess would be A or B

OpenStudy (math&ing001):

It's B

OpenStudy (anonymous):

Thank you

OpenStudy (math&ing001):

You're welcome

OpenStudy (anonymous):

Can you help me with this one ? A sequence has its first term equal to 4 and each term of the sequence is obtained by adding 2 to the previous term. If f(n) represents the nth term of the sequence, which of the following recursive functions best defines this sequence? f(1) = 2 and f(n) = f(n - 1) + 4; n > 1 f(1) = 4 and f(n) = f(n - 1) + 2n; n > 1 f(1) = 2 and f(n) = f(n - 1) + 4n; n > 1 f(1) = 4 and f(n) = f(n - 1) + 2; n > 1

OpenStudy (anonymous):

both please

OpenStudy (math&ing001):

What's the first term of your sequence?

OpenStudy (anonymous):

= to 4 right ?

OpenStudy (math&ing001):

Anyway the answer is f(1) = 4 and f(n) = f(n - 1) + 2; n > 1

OpenStudy (jdoe0001):

say .... we know that the 1st term is 4 so that pretty much rules out 2 options

OpenStudy (jdoe0001):

the \(\bf n^{th}\) term, that is the f(n), is whatever the previous term was, PLUS 2

OpenStudy (jdoe0001):

let's say..... if my \(\bf n^{th}\) term is say .... 5.... what would be the previous term order?

OpenStudy (anonymous):

For anyone looking for the correct answer for the original question: If f(x) is the height, in cm, of a sunflower plant which is x days old, which of the following statements best describes the meaning of f(60) = 210 The answer is A because f(x)= the height in cm which is 60 X= how many days old the flower is which is 210 I take FLVS and just took this tests and got it correct don't listen to that dude he doesn't know what e is talking about.

OpenStudy (anonymous):

thank you ! @akeller27